what equation has a line that passes through (4,6) and has a slope of 1/2
Answer by macston(5194)

m=1/2 Point=(4,6)
y=mx+b
To calculate b we use the x and y of the point
6=1/2(4)+b
6=2+b Subtract 2 from each side.
4=b
ANSWER: The equation in slope-intercept form is: y=(1/2)x+4.